Circular No. 7027

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1946+274 R. M. Hjellming, National Radio Astronomy Observatory (NRAO);

and A. J. Mioduszewski, Joint Institute for VLBI in Europe and NRAO, report: “Very Large Array observations on Oct. 7.02 UT reveal a 5-mJy radio source at 1.49 GHz inside the error region for the transient x-ray pulsar XTE J1946+274 reported by Takeshima et al. (IAUC 7016; see also Smith et al., IAUC 7014). This radio source has a position α = 19^{h}45^{m}31^{s}.75, δ = +27^{o}21′38′′.0 (equinox 2000.0; 2′′ uncertainty). Since it was at < 1 mJy at 4.9 GHz on Oct. 7.02, and it was not detected at 1.49 GHz with upper limits of 1 and 2 mJy on Sept. 21 and 27, respectively, this may indicate that we have detected the early rise of an optically thick radio source. The radio source was also not present in a 1995 May 16 NVSS survey image at 1.4 GHz with a limit of ∼ 2 mJy. Further radio observations are needed to confirm variability, particularly since it is unusual for a transient x-ray pulsar to exhibit detectable radio continuum emission.”

NGC 7582 J. P. Halpern, L. E. Kay, and K. M. Leighly, Columbia University,

write: “We obtained spectropolarimetry (range 390–750 nm; 0.3-nm resolution) of the Seyfert-2 galaxy NGC 7582 on June 20 using the Cerro Tololo 4-m telescope. In a 24-min exposure, we do not see the broad Balmer lines detected by Joguet et al. (IAUC 7024) on July 11, in either direct or polarized flux. If the emergence of this transient broad-line emission was caused either by the removal of absorbing material at a velocity of 25 000 km/s (the reported FWZI of the lines), or by the tidal disruption of a star with such a relative velocity, the 21-day interval sets an upper limit of 4.5 × 10^{15} cm on the size of the emission-line region on July 11.”

DRACONID METEORS 1998 Numerous reports have been received of meteor activity coming from

the Draconid/Giacobinid radiant, with a visual peak of around 40–50 meteors/hr seen around Oct. 8.56 UT (S. Yoshida and D. Ito in Japan). Peaks near 100/hr were detected with a high-density video camera around Oct. 8.5 by J. Watanabe at the Dodaira Observatory, and a similar peak time and rates were detected via forward-scattered radio echoes by K. Suzuki (Aichi, Japan) and by E. P. Bus (Groningen, The Netherlands). Backscattering radar echoes by J. Borovička, P. Prida, and P. Pecina (Ondřejov Observatory) showed high activity perhaps peaking as late as Oct. 8.6.

1998 October 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green